A man has a collection of dimes and quarters with a total of $3.50. If he has 7 more dimes than quarters, how many of each coins does he have?
Found 2 solutions by math_helper, Alan3354:
Answer by math_helper(2461)


q = number of quarters
d = number of dimes = q+7

25q + 10(q+7) = 350 (cents)
35q + 70 = 350
35q = 280
q = 280/35 = 8 ---> d = 15
=====
Ans: 8 quarters and 15 dimes
=====
Check:
25(8) + 10(15) = 200 + 150 = 350 (= $3.50) (ok)

Answer by Alan3354(69443)

A man has a collection of dimes and quarters with a total of $3.50. If he has 7 more dimes than quarters, how many of each coins does he have?
======================
Set the "7 more dimes" aside.
Now it's pairs of 1Q & 1D worth $2.80
280/35 = 8 pairs.
+ 7 dimes ---> 8Q & 15D